OpenStudy (anonymous):

I would like to know what is the next line if you are showing your work for the problem X^2 +2x-8=0 solve for X

OpenStudy (radar):

\[X ^{2}+2X-8=0 FACTORING: (X-4)(X+2=0\]

OpenStudy (radar):

X-4=0 OR X=4 X+2=0 OR X=-2